CLASS INFORMATION AND POLICY

COURSE: Calculus III, Math208 Sec 5, Fall 2001

TEXT: Calculus, Single and Multivariable, , by Hughes-Hallet, et al, 2nd ed., John Wiley

& Sons, Inc. (1998).

Calculator: A graphing calculator, TI-85 or TI-86, is recommended. TIME: 11:30—12:20 pm MTWF ROOM: M&N 206

INSTRUCTOR: Dr. Bo Deng

OFFICE: 828 Oldfather Hall

OFFICE PHONE: 472-7219(no message please)

E-MAIL ADDRESS: bdeng@math.unl.edu

OFFICE HOURS: 9:30—10:20 pm M W F(subject to change) or by appointment

Attendance: A semester total of 6 non-excusable absence will result in a half letter grade deduction. A semester total of 10 class absence will result in a full letter grade deduction. A perfect attendance record receives 10 bonus points.

Quiz: Quiz every Monday on homework type problems on sections covered in the previous week, except holidays or exams scheduled on the same weeks. All quizzes together count no more than 100 points towards your course grade.

Exams: There will be three hour exams plus the final exam. The hour exams will consist of both routine problems similar to your homework assignment and problems which may require some thought. The final exam is a uniform, comprehensive exam. Each hour exam counts 100 points towards your course grade. The final exam counts 200 points.

Group Projects: There will be no more than two group projects this semester. Each group should consist of two members: you and your project partner. Only in some rare situations will a three-member-group be allowed. Your group assignment will be based on your academic standing prior to the assignment. Your group must work independently from others. Shared presentation materials in any form and to any extent with others will not be accepted, and all parties involved in violating this rule will be penalized equally. Late projects will not be accepted. More information will be forthcoming when the projects are assigned. Each project counts no more than 20 points towards your course grade.

Web Tests: You must take two Web Tests as described in your Syllabus. If you get at least 8 questions out of 10 right, you get at least 80%, but if you get fewer right your score will be below 50%, so the grades on those tests are A, B, F. Each web test count counts no more than 20 points towards your course grade.

Grade: Throughout the semester you must earn credits from the quizzes, group projects, web tests, hour exams, and the final exam against a grand total of no more than 700 points. The standard letter grade will be assigned to your course grade based on your total point against the grand total.

A Few Suggestions: Success in any mathematics course requires a tendency to perfectionism. Every step of the way, strict attention to the smallest detail is absolutely necessary. If you can learn to acquire the trait of perfectionism for certain courses, you will succeed much more easily in mathematics. The other way around is even truer, that is, if you are good at learning mathematics you are probably also good at learning other subjects as well.

It is important for you to work out the homework problems as much independently as possible. Try to reason through problems even though there are formulas or algorithms ready to use. This reasoning process is the essential part of the mathematical thought process. You should also do your homework in a timely fashion. Like most math courses, the material is very cumulative and therefore is easy to get behind. Time can be your friend as much as your foe.

Always read ahead. I will plan class activities assuming you have done the reading. As you read the text, concentrate on the general development first. Details of calculus and algebra are generally left to the reader; and eventually the reader should confirm all calculations. As a general rule, one should consider a pencil and scratch paper as essential equipment for reading mathematics.

Should you have any questions, speak up in class, seek individual help from me.

End Page